OpenStudy (anonymous):

2.5 qt/min = ____ gal/h

OpenStudy (austinl):

For this one I would suggest dimensional analysis. Are you familiar with it?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

this is for ratios and rates

OpenStudy (austinl):

I know, I will set it up and you can solve. \(\dfrac{2.5qt}{1min}\times\dfrac{??min}{1hour}\times\dfrac{1gal}{??qt}=\dfrac{??gal}{hr}\) How many minutes in one hour? How many quarts in one gallon? Then you just multiply!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

there are 60 min in an hr

OpenStudy (austinl):

Correct, and how many quarts in one gallon?

OpenStudy (anonymous):

4 quarts

OpenStudy (austinl):

Correctomundo! So just plug those in and multiply across top and bottom!

OpenStudy (anonymous):

yo speak spanish right???

OpenStudy (anonymous):

so multiply \[1/4\]

OpenStudy (anonymous):

sorry multiply 60 x 4 and 1x 1 right???

OpenStudy (anonymous):

60/1 and 1/4

OpenStudy (austinl):

I can speak it a little, haven't really spoken it in over 2 years. \(\dfrac{2.5qt}{1min}\times\dfrac{60min}{1hr}\times\dfrac{1gal}{4qt}=\dfrac{2.5\times60\times1}{1\times1\times4}=~?\)

OpenStudy (anonymous):

i got \[150 gal /4 hr\]

OpenStudy (austinl):

Which when divided is? 150/4?

OpenStudy (anonymous):

37.5

OpenStudy (austinl):

Correct!

OpenStudy (anonymous):

:) pat on the shoulder hahha lol